Subshilava:

f(x) = (x + 5) * (x + 3) ^ 2 * (x + 2) ^ 2 What is the error in the equation, if any. The exponent on the (x + 5) factor should be even. no error The first factor should be (x - 5) rather than (x + 5) The exponent on the (x + 3) factor should be odd. The last factor should be (x - 2) rather than (x + 2)

XShawtyX:

The exponent on the (x + 5) factor should be even. The given equation is 𝑓(𝑥)=(𝑥+5)(𝑥+3)2(𝑥+2)2f(x)=(x+5)(x+3) 2(x+2) 2 , and the only issue stated is that the exponent of (𝑥+5)(x+5) should be even. Since it's currently raised to the power of 1 which is odd), this is the error that needs to be corrected.

Subshilava:

@xshawtyx wrote:
The exponent on the (x + 5) factor should be even. The given equation is 𝑓(𝑥)=(𝑥+5)(𝑥+3)2(𝑥+2)2f(x)=(x+5)(x+3) 2(x+2) 2 , and the only issue stated is that the exponent of (𝑥+5)(x+5) should be even. Since it's currently raised to the power of 1 which is odd), this is the error that needs to be corrected.
How can you tell what errors are in a problem?

XShawtyX:

So uh to Indentify Errors you compare the provided equation with any specific rules or conditions given in the problem.

XShawtyX:

I can break it down step by step On how I did it If u still don't understand